OpenStudy (anonymous):

Which is an equation in slope-intercept form for the line passing through (–1, 5) and (5, 8)

OpenStudy (imstuck):

First find the slope using the slope-intercept formula. That is\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]Filling in these using our two points, we get \[\frac{ 8-5 }{ 5-(-1)}=\frac{ 3 }{ 6 }=\frac{ 1 }{ 2 }\]So the slope is 1/2. Now we use the point-slope formula to find the equation. \[y-y _{1}=m(x-x _{1})\]Pick whichever point you would like to use here. Lets use the one with no negatives in it. Your y1 is your y-coordinate of that point and your x1 is your x coordinate of that point, namely y = 8 and x = 5. So let's do this.\[y-8=1/2(x-5)\]Notice that we subbed in the slope of 1/2. Now distribute the 1/2 into the parenthesis to get \[y-8=\frac{ 1 }{ 2 }x-\frac{ 5 }{ 2 }\]Move the -8 to the other side using addition and the fact that 8 is the same thing as 16/2.\[y=\frac{ 1 }{ 2 }x-\frac{ 5 }{ 2 }+\frac{ 16 }{ 2 }\]This all simplifies down to \[y=\frac{ 1 }{ 2 }x+\frac{ 11 }{ 2 }\]And that's it in a nutshell. And goodnight for now. I'll return in the morning!